OpenStudy (anonymous):

Find the polar coordinates of the point whose rectangular coordinates are (-3,-3). a.) (-3 sqrt(2),5 pi /4) b.) (3 sqrt(2),pi /4) c.) (18,pi /4) d.) (3 sqrt(2),5 pi /4)

OpenStudy (superdavesuper):

to convert (x,y) to (r, theata) use the eqns: x=r*sin(theata) and y=r*cos(theata)

OpenStudy (anonymous):

You need to find r and \( \theta\). In which quadrant is (-3,-3} and what is the distance to the origin?

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

How do you find the distance from (0,0) to (-3,-3)? \[ r=\sqrt{( -3-0)^2+(-3-0)^2}=\sqrt{18}=\sqrt{ 9(2)}= 3 \sqrt 2 \]

OpenStudy (anonymous):

We have to chose between b and d, using the angle \(\theta\) Can you do that?

OpenStudy (anonymous):

distance is 3 sqrt(2) and angle is pi/4.

OpenStudy (anonymous):

The angle pi/4 will put you in the first quadrant and your point is in the third quadrant

OpenStudy (anonymous):

r is 3sqrt(2) and theata is 5pi/4.....d right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yay! thanks so much.

OpenStudy (anonymous):

YW